b. Prove that if a is an odd integer, then a | b2 – 1 implies that a = (a, b – 1)(a, b + 1).
Answer #1

Answedia Given data TO prove that if a is an Odd integer, then a/6²_11 implies that a = (ab-1) (0,b)) pdoof. Suppose a dividea= pq | (6.1). To apply ged (6-1, a) -(6-1, pd) - p Similarly (b+10) (btl, PP. 1-4 so a=pd = (a,b-1).(0,6+1). (a,b-1) (Q,b+1)